OpenStudy (anonymous):

What is the first step to solve the following equation?x^2-2x=0 A.use the distributive property B.divide by x C.use the zero product property D.factor out x Please Help

OpenStudy (solomonzelman):

well, factoring out of x is the easiest to me .

OpenStudy (solomonzelman):

x²-2x=0 x(x-2)=0 x=0 or x-2=0 x=2

OpenStudy (anonymous):

Ok thanks! I was thinking either that or divide by x

OpenStudy (solomonzelman):

Well, dividing by x, wouldn't give the solution of x=0. It would just cancel the x.